OpenStudy (omgthatgirl):

Simplify the imaginary number square root -17

OpenStudy (jhannybean):

\[\large \sqrt{-17} = \sqrt{-1}\cdot \sqrt{17}\]\[\large i = \sqrt{-1}\]Now substitute and simplify.

OpenStudy (omgthatgirl):

Ok would it be 0i-17?

OpenStudy (omgthatgirl):

@Jhannybean

OpenStudy (jhannybean):

what is 0i mean?

OpenStudy (omgthatgirl):

lol i don't know

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\large \sqrt{-17} = \color{blue}{\sqrt{-1}}\cdot \sqrt{17}\]\[\large \color{blue}{i = \sqrt{-1}}\]Now substitute and simplify. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (jhannybean):

Replace the blue part.

OpenStudy (omgthatgirl):

blue part? @Jhannybean

OpenStudy (jhannybean):

Look at my above post.